# Shaping Sparse Rewards in Reinforcement Learning: A Semi-supervised Approach

### v1

@@ A.1 The Gradient of $L_{QV}$
$L_{QV}=\begin{cases}0,&Q(s_{t},a_{t},\theta_{1})-V(s_{t},\theta_{2})<0,\\ \displaystyle\frac{1}{\mu B}\sum_{t=1}^{\mu B}\left(Q(s_{t},a_{t},\theta_{1})-V(s_{t},\theta_{2})\right)^{2},&\text{else}.\end{cases}$
@@
Denote $\delta_{t}(\theta)=Q(s_{t},a_{t},\theta_{1})-V(s_{t},\theta_{2})$. Consider a batch of size $B$, within which a proportion $\mu_{l}\geq 0$ of the samples satisfy $\delta_{t}(\theta)<0$. The remaining proportion $(1-\mu_{l})$ contributes to the gradient with respect to the parameter $\theta=(\theta_{1},\theta_{2})$ is given by:
@@
$\frac{\partial L_{QV}}{\partial\theta}=\frac{2}{(1-\mu_{l})B}\sum_{t\in\mathcal{I}_{+}}Q(s_{t},a_{t},\theta_{1})-V(s_{t},\theta_{2})\cdot\begin{bmatrix}\nabla_{\theta_{1}}Q(s_{t},a_{t},\theta_{1})\\ -\nabla_{\theta_{2}}V(s_{t},\theta_{2})\end{bmatrix}.$
@@
A.2 The Gradient of $L_{r}$
Consider loss function $L_{r}$ in Eq.(4) as follow
$L_{r}=\frac{1}{\mu B}\sum_{t=1}^{\mu B}\mathbbm{1}(\max(\mathbf{q_{t}})\geq\lambda)\left(r_{t}-\alpha(\mathbf{q_{t}},\lambda)\right)^{2},$
where confidence score $\mathbf{q_{t}}=\beta Q(s_{t},a_{t},\theta_{1})+(1-\beta)V(s_{t+1},\theta_{2})$. Note that the indicator function $\mathbbm{1}(\max(\mathbf{q_{t}})\geq\lambda)$ is not differentiable due to its discontinuity at the boundary of $\max(\mathbf{q_{t}})\geq\lambda$. The derivative of the indicator function only has non-zero contribution at the boundary, i.e. $\max(\mathbf{q_{t}})=\lambda$. In order to derive the gradient of $L_{r}$, we approximate the indicator function $\mathbbm{1}(\max(\mathbf{q_{t}})\geq\lambda)$ using differentiable function, for simplicity, the sigmoid function, achieving the effect of approximate binary selection:
$\mathbbm{1}(\max(\mathbf{q_{t}})\geq\lambda)\approx\sigma(\max(\mathbf{q_{t}})-\lambda),$
@@ A.2 The Gradient of $L_{r}$
Since $\frac{\partial\sigma(x)}{\partial x}=\sigma(x)(1-\sigma(x))$, plugging it in Eq.(7) we get:
$\frac{\partial f_{r}}{\partial\theta_{1}}=\sigma(\max(\mathbf{q_{t}})-\lambda)(1-\sigma(\max(\mathbf{q_{t}})-\lambda))\cdot\frac{\partial\max(\mathbf{q_{t}})}{\partial\mathbf{q_{t}}}\cdot\beta\frac{\partial Q(s_{t},a_{t},\theta_{1})}{\partial\theta_{1}}.$
Now, the gradient of $L_{r}$with respect to $\theta_{1}$ and $\theta_{2}$ can be computed using the chain rule. The gradient with respect to $\theta_{1}$ is:
@@ A.3 The Gradient of $L_{s}$
$L_{s}=\frac{1}{(1-\mu)B}\sum_{t=1}^{(1-\mu)B}\left[\mathbbm{1}(\max(\mathbf{q_{t}^{s}})\geq\lambda,\,\max(\mathbf{q_{t}^{w}})\geq\lambda)\times\mathcal{H}\left(\mathbf{q_{t}^{w,\,\prime}},\,\mathbf{q_{t}^{s,\,\prime}}\right)\right].$
To compute $\partial L_{s}/\partial\theta_{1}$, we will apply the chain rule. The loss $L_{s}$ involves a sum of products, and we need to consider the parts that depend on $\theta_{1}$. We adopt the same approach as in Section A.2, using sigmoid function (see Eq.(6)) approximation for the indicator function. Therefore, the overall approximation of the indicator function is:
$\mathbbm{1}(\max(\mathbf{q_{t}^{s}})\geq\lambda,\max(\mathbf{q_{t}^{w}})\geq\lambda)\approx\sigma(\max(\mathbf{q_{t}^{s}})-\lambda)\cdot\sigma(\max(\mathbf{q_{t}^{w}})-\lambda).$ (9)
Similar to that in Appendix.A.2, the approximate derivative of the indicator function $\mathbbm{1}(\max(\mathbf{q_{t}^{s}})\geq\lambda,\,\max(\mathbf{q_{t}^{w}})\geq\lambda)$ (denoted as $f_{s}$) with respect to $\theta_{1}$ is computed based on Eq.(8) and Eq.(9):
$\frac{\partial f_{s}}{\partial\theta_{1}}=$ $\sigma(\max(\mathbf{q_{t}^{s}})-\lambda)(1-\sigma(\max(\mathbf{q_{t}^{s}})-\lambda))\cdot\sigma(\max(\mathbf{q_{t}^{w}})-\lambda)\cdot\frac{\partial\max(\mathbf{q_{t}^{s}})}{\partial\theta_{1}}$
$+\sigma(\max(\mathbf{q_{t}^{s}})-\lambda)\cdot\sigma(\max(\mathbf{q_{t}^{w}})-\lambda)(1-\sigma(\max(\mathbf{q_{t}^{w}})-\lambda))\cdot\frac{\partial\max(\mathbf{q_{t}^{w}})}{\partial\theta_{1}}.$ (10)
@@
The loss $L_{s}$ involves a sum of products, and now that we have derived the parts that depend on $\theta_{1}$ in Eq.(A.3) and Eq.(11), the gradient of the loss function with respect to $\theta_{1}$ is:
$\frac{\partial L_{s}}{\partial\theta_{1}}=\frac{1}{(1-\mu)B}\sum_{t=1}^{(1-\mu)B}\left[\frac{\partial f_{s}}{\partial\theta_{1}}\cdot\mathcal{H}\left(\mathbf{q_{t}^{w,\,\prime}},\,\mathbf{q_{t}^{s,\,\prime}}\right)-f_{s}\cdot\frac{1}{q_{i^{*}}}\cdot\frac{\partial q_{i^{*}}}{\partial\theta_{1}}\right].$
where
$\frac{\partial f_{s}}{\partial\theta_{1}}=\sigma(\max(\mathbf{q_{t}^{s}})-\lambda)(1-\sigma(\max(\mathbf{q_{t}^{s}})-\lambda))\cdot\sigma(\max(\mathbf{q_{t}^{w}})-\lambda)\cdot\frac{\partial\max(\mathbf{q_{t}^{s}})}{\partial\theta_{1}}$
$+\sigma(\max(\mathbf{q_{t}^{s}})-\lambda)\cdot\sigma(\max(\mathbf{q_{t}^{w}})-\lambda)(1-\sigma(\max(\mathbf{q_{t}^{w}})-\lambda))\cdot\frac{\partial\max(\mathbf{q_{t}^{w}})}{\partial\theta_{1}}.$
@@
Appendix B Supplementary Materials for Experiments
B.1 Hyperparameters Settings
The hyperparameter setup of our experiments in Section 4 consists of the data augmentation parameters, the SSRS framework parameters and other general configurations. As for data augmentation parameters, we adopt consistent settings throughout the experiments in Section 4.1-4.3. Due to different environments, the shapes of states and observations are different, so the parameter $n$ of double entropy data augmentation and cutout data augmentation, which is related to the state and observation dimension, are different in Atari and Robotic environments (see Table 2).
Value of $n$ Double-Ent Cutout
Atari 8 16
Robotic 4 3
In Table 3, we present the settings of hyperparameters in SSRS framework, which are based on preliminary hyperparameter tuning results as we mentioned in Section 4. Note that dynamic setting is applied to the consistency coefficient $\alpha$ and the threshold of confidence $\lambda$, i.e., the value of these hyperparameters change during the training process, and the value shown in the Table 3 is the final value of the dynamic setting. In detail, the confidence threshold $\lambda$ grows along with the training iteration. In a training process with $N_{iter}$ iteration in total, $\lambda$ is as followed at the $n_{iter}$ iteration:
$\lambda=0.6+0.3\times\left(1-e^{-\frac{n_{iter}}{N_{iter}}}\right)$
Similarly, in a training process with $N_{iter}$ iteration in total, the consistency coefficient $\alpha$ is as followed at the $n_{iter}$ iteration:
$\alpha=\begin{cases}0.2+(0.7-0.2)\times\left(\frac{n_{iter}}{N_{iter}\times 0.8}\right),&\text{if }n_{iter}<0.8\cdot N_{iter}\\ 0.7,&\text{otherwise}\end{cases}$
Table 3: The setting of hyperparameters in SSRS framework.
hyperparameter description value
$\beta$ the QV confidence coefficient 0.5
$\lambda$ the threshold of confidence 0.9
$\alpha$ the consistency coefficient 0.7
$p_{u}$ the update probability 0.01
$N_{z}$ the size of the fixed reward set $Z$ 12
$n$ partition number of Double Entropy Augmentation 8
The network architecture, training parameters, and other hyperparameters used in the experiments for the base DQN, DDPG, PER and ICM algorithms follow the settings provided by the Tianshou RL Library Weng u. a. (2022) and each paper Mnih u. a. (2015); Lillicrap u. a. (2015); Schaul u. a. (2015); Pathak u. a. (2017). The hyperparameter settings of RCP refer to the settings in the code base Bricken (2021) and paper Kumar u. a. (2019). Additionally, the network architecture of the reward estimator is also presented in Table 4.
Table 4: Network architecture and hyperparameters
architecture and hyperparameters Value
Value Network - Conv(32-8x8-4) / ReLU
- Conv(64-4x4-2) / ReLU
- Conv(64-3x3-1) / ReLU
- Flatten
- FC(512) / ReLU
- FC(|A|)
Reward - FC(128) / ReLU
- Dropout(p=0.2)
- FC(64) / ReLU
- Dropout(p=0.2)
- FC(32) / ReLU
- Dropout(p=0.2)
- FC($N_{z}$) / ReLU
- Softmax(dim=-1)
DQN Discount Factor $\gamma$ 0.99
Learning Rate 0.0001
Batch Size 256
Number of Parallel Environments 8
Estimation Step 3
Target Update Frequency 500
DDPG Discount Factor $\gamma$ 0.99
Actor Learning Rate 0.001
Critic Learning Rate 0.001
Batch Size 256
Number of Parallel Environments 8
Update Per Step 1
Soft Update $\tau$ 0.005
Exploration Noise 0.1
PER Priority bias $\alpha$ 0.6
Importance Sampling correction $\beta$ 0.4
ICM ICM Learning Rate Scale 0.2
Reward Scale 0.01
Forward Loss Scale 0.2
Intrinsic Curiosity Moudule Hidden Size=[512]
RCP $\beta$ Reward Weighting 1
Max Loss Weighting 20
TD $\lambda$ 0.95
B.2 List of Data Augmentation Methods
Various types of data augmentation methods are used in the process of consistency regularization. Since the transformation double entropy data augmentation is thoroughly stated in Section 3.1, we here list all transformation operations for data augmentation strategies involved in our experiment in Table 5 for completeness. The parameter setup of the transformation refers to the Appendix B.1.
Regarding to the heterogeneous SSRS variants which apply two weak and strong data augmentations in consistency regularization, the transformation details of the legends in Figures 5 are listed below:
  • SSRS-S: SSRS-S applies $\sigma$ random gauss noise as the weak augmentation and double entropy data augmentation as strong augmentation.
  • SSRS-M: SSRS-M applies $\sigma$ random gauss noise as the weak augmentation and smooth data augmentation as strong augmentation.
  • SSRS-C: SSRS-C applies $\sigma$ random gauss noise as the weak augmentation and cutout data augmentation as strong augmentation.
Table 5: Data augmentation methods in the control group and SSRS variants. Weak augmentation is gauss augmentation by default, and the rest augmentations are used as strong augmentations respectively.
Transformation Description Parameter Range
gauss add $\sigma$ random gauss noise to the observation $\sigma$ $default=0.1$
cutout randomly zero out $n$ columns of observation $n$ $default=16$
smooth smooth $n$ consecutive observations $n$ $default=3$
scale scale the observation by random scale factor $\lambda$ $\lambda$ $(0.8,1.2)$
translation horizontally translate the observation by factor $\lambda$ $\lambda$ $(0,0.1)$
flip horizontally mirror flip the observation

### v2

@@ Appendix C Hyperparameters Settings
$\alpha=\begin{cases}0.2+(0.7-0.2)\times\left(\frac{t}{T\times 0.8}\right),&\text{if }t<0.8\cdot T\\ 0.7,&\text{otherwise}\end{cases}$
Table 4: The setting of hyperparameters in SSRS framework.
hyperparameter description value
$\beta$ the QV confidence coefficient 0.5
$\lambda$ the threshold of confidence 0.9
$\alpha$ the consistency coefficient 0.7
$p_{u}$ the update probability 0.01
$N_{z}$ the size of the fixed reward set $Z$ 12
$n$ partition number of Double Entropy Augmentation 8
The network architecture, training parameters, and other hyperparameters used in the experiments for the base SAC and DDPG algorithms follow the settings provided by the Tianshou RL Library (Weng et al. 2022) and paper Haarnoja et al. (2019); Lillicrap et al. (2015). Note the critic network shares the same hidden size with actor network and we only detail the actor network in Table 5. The hyperparameter settings of RCP refer to the settings in the code base (Bricken 2021) and paper (Kumar, Peng, and Levine 2019). The hyperparameter settings of SORS refer to the settings in the code base (Hiwonjoon 2021) and paper (Memarian et al. 2021). Additionally, the network architecture of the reward estimator is also presented in Table 5.
Table 5: Network architecture and hyperparameters
architecture and hyperparameters Value
Actor Network - Conv(32-8x8-4) / ReLU
- Conv(64-4x4-2) / ReLU
- Conv(64-3x3-1) / ReLU
- Flatten
- FC(512) / ReLU
- FC(—A—)
Reward Network - FC(128) / ReLU
- Dropout(p=0.2)
- FC(64) / ReLU
- Dropout(p=0.2)
- FC(32) / ReLU
- Dropout(p=0.2)
[…]
- Softmax(dim=-1)
DDPG Discount Factor $\gamma$ 0.99
Actor Learning Rate 0.001
Critic Learning Rate 0.001
Batch Size 256
Number of Parallel Environments 8
Update Per Step 1
Soft Update $\tau$ 0.005
Exploration Noise 0.1
SAC Entropy Coefficient $\alpha$ 0.05
Auto-Alpha True
Alpha Learning Rate 0.0003
RCP $\beta$ Reward Weighting 1
Max Loss Weighting 20
TD $\lambda$ 0.95
SORS Reward Update Period 1000
Reward Update Num 100
Reward Batch Size 10
Appendix D The Gradients of Loss Functions
The Gradient of $L_{QV}$
The $L_{QV}$ has the following explicit form:
@@ Appendix D The Gradients of Loss Functions
Denote $\delta_{t}(\theta)=Q(s_{t},a_{t},\theta_{1})-V(s_{t},\theta_{2})$, where $\theta$ is the combination of $\theta_{1}$ and $\theta_{2}$. Consider a batch of size $B$, within which a proportion $\mu_{l}\geq 0$ of the samples satisfy $\delta_{t}(\theta)<0$. The remaining proportion $(1-\mu_{l})$ contributes to the gradient with respect to the parameter $\theta=(\theta_{1},\theta_{2})$ is given by:
@@
$\begin{split}\frac{\partial L_{QV}}{\partial\theta}=\frac{2}{(1-\mu_{l})B}\sum_{t\in\mathcal{I}_{+}}&\Bigl{(}Q(s_{t},a_{t},\theta_{1})-V(s_{t},\theta_{2})\Bigr{)}\\ &\cdot\begin{bmatrix}\nabla_{\theta_{1}}Q(s_{t},a_{t},\theta_{1})\\ -\nabla_{\theta_{2}}V(s_{t},\theta_{2})\end{bmatrix}.\end{split}$
@@
The Gradient of $L_{r}$
Consider loss function $L_{r}$ in Eq.(3) as follow:
$L_{r}=\frac{1}{\mu B}\sum_{t=1}^{\mu B}\mathds{1}(\max(\mathbf{q_{t}})\geq\lambda)\left(r_{t}-\alpha(\mathbf{q_{t}},\lambda)\right)^{2},$
where confidence score $\mathbf{q_{t}}=\beta Q(s_{t},a_{t},\theta_{1})+(1-\beta)V(s_{t+1},\theta_{2})$. Note that the indicator function $\mathds{1}(\max(\mathbf{q_{t}})\geq\lambda)$ is not differentiable due to its discontinuity at the boundary of $\max(\mathbf{q_{t}})\geq\lambda$. The derivative of the indicator function only has non-zero contribution at the boundary, i.e. $\max(\mathbf{q_{t}})=\lambda$. In order to derive the gradient of $L_{r}$, we approximate the indicator function $\mathds{1}(\max(\mathbf{q_{t}})\geq\lambda)$ using differentiable function, for simplicity, the sigmoid function, achieving the effect of approximate binary selection:
$\mathds{1}(\max(\mathbf{q_{t}})\geq\lambda)\approx\sigma(\max(\mathbf{q_{t}})-\lambda),$
@@
Since $\frac{\partial\sigma(x)}{\partial x}=\sigma(x)(1-\sigma(x))$, plugging it in Eq.(9), we get:
$\frac{\partial f_{r}}{\partial\theta_{1}}=\sigma(\max(\mathbf{q_{t}})-\lambda)(1-\sigma(\max(\mathbf{q_{t}})-\lambda))\\ \cdot\frac{\partial\max(\mathbf{q_{t}})}{\partial\mathbf{q_{t}}}\cdot\beta\frac{\partial Q(s_{t},a_{t},\theta_{1})}{\partial\theta_{1}}.$
Similarly, we can also derive $frac{\partial f_{r}}{\partial\theta_{2}}$. Now, the gradient of $L_{r}$with respect to $\theta_{1}$ and $\theta_{2}$ can be computed using the chain rule. The gradient with respect to $\theta_{1}$ is:
@@
$L_{s}=\frac{1}{(1-\mu)B}\sum_{t=1}^{(1-\mu)B}\begin{aligned} &\left[\mathds{1}(\max(\mathbf{q_{t}^{s}})\geq\lambda,\,\max(\mathbf{q_{t}^{w}})\geq\lambda)\right.\\ &\left.\times\mathcal{H}\left(\mathbf{q_{t}^{w,\,\prime}},\,\mathbf{q_{t}^{s,\,\prime}}\right)\right].\end{aligned}$
To compute $\partial L_{s}/\partial\theta_{1}$, we will apply the chain rule. The loss $L_{s}$ involves a sum of products, and we need to consider the parts that depend on $\theta_{1}$. We adopt the same approach, using sigmoid function (see Eq.(8)) approximation for the indicator function. Therefore, the overall approximation of the indicator function is:
$\mathds{1}(\max(\mathbf{q_{t}^{s}})\geq\lambda,\max(\mathbf{q_{t}^{w}})\geq\lambda)$ (11)
$\approx$ $\sigma(\max(\mathbf{q_{t}^{s}})-\lambda)\cdot\sigma(\max(\mathbf{q_{t}^{w}})-\lambda).$
The approximate derivative of the indicator function $\mathds{1}(\max(\mathbf{q_{t}^{s}})\geq\lambda,\,\max(\mathbf{q_{t}^{w}})\geq\lambda)$ (denoted as $f_{s}$) with respect to $\theta_{1}$ is computed based on Eq.(10) and Eq.(11):
$\begin{split}\frac{\partial f_{s}}{\partial\theta_{1}}={}&\sigma(\max(\mathbf{q_{t}^{s}})-\lambda)\bigl{(}1-\sigma(\max(\mathbf{q_{t}^{s}})-\lambda)\bigr{)}\\ &\cdot\sigma(\max(\mathbf{q_{t}^{w}})-\lambda)\cdot\frac{\partial\max(\mathbf{q_{t}^{s}})}{\partial\theta_{1}}\\ +{}&\sigma(\max(\mathbf{q_{t}^{s}})-\lambda)\cdot\sigma(\max(\mathbf{q_{t}^{w}})-\lambda)\\ &\cdot\bigl{(}1-\sigma(\max(\mathbf{q_{t}^{w}})-\lambda)\bigr{)}\cdot\frac{\partial\max(\mathbf{q_{t}^{w}})}{\partial\theta_{1}}.\end{split}$ (12)
@@
The loss $L_{s}$ involves a sum of products, and now that we have derived the parts that depend on $\theta_{1}$ in Eq.(12) and Eq.(13), the gradient of the loss function with respect to $\theta_{1}$ is:
$\frac{\partial L_{s}}{\partial\theta_{1}}=\frac{1}{(1-\mu)B}\sum_{t=1}^{(1-\mu)B}\begin{aligned} &\left[\frac{\partial f_{s}}{\partial\theta_{1}}\cdot\mathcal{H}\left(\mathbf{q_{t}^{w,\,\prime}},\,\mathbf{q_{t}^{s,\,\prime}}\right)\right.\\ &\left.-f_{s}\cdot\frac{1}{q_{i^{*}}}\cdot\frac{\partial q_{i^{*}}}{\partial\theta_{1}}\right].\end{aligned}$
where $\frac{\partial f_{s}}{\partial\theta_{1}}$ refers to Eq.(12).